Write first five multiples of: 
9 ??????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The required multiples are:
9 ×1 =9
9 × 2 = 18
9 × 3 = 27
9 × 4 = 36
9 × 5 = 45
Hence, the first five multiples of 9 are 9, 18, 27, 36 and 45

Thanks